OpenStudy (dil.howlter):

Keisha is reading a 350 page book at a rate of 27 pages per day. Use a point-slope equation to determine how many pages she will have left to read after reading the book for 10 days. How many pages will Keisha have left to read after 10 days? Lizzy is tiling a kitchen floor for the first time. She had a tough time at first and placed only 9 tiles the first day. She started to go faster and by the end of day 4, she had placed 39 tiles. She worked at a steady rate after the first day. Use an equation in point slope form to determine how many days Lizzy took to place all of the 100 tiles

OpenStudy (zpupster):

Hint watch your units if slope(m) is y/x it will be a unit / another unit so y - y1=m(x-x1) we know x1(starting days) is 0 and we know y1 is 350( end of book) plugging in y-350=27(10-0)

OpenStudy (zpupster):

in the case of Lizzy day 1 she placed 5 tiles ( x1,y1) day 4 " " 35 ( x2,y2) tiles/day= y2-y1/x2-x1= m = slope plugging in 35-5/4-1 solve and this will give you m or our slope Next: 100 is y 5 is y1 x=how many days x1 = 1 plugging in y - y1=m(x-x1) 100-5=m(x-1) plug in m from abaove solve for x and that will give total amount of days